OpenStudy (anonymous):

a) What number is 1/4 of the way from 16 to 256? b) What common fraction is one-third of the way from 1/4 to 1/3 ? Please show full work! thank you

OpenStudy (anonymous):

256-16=240 (1/4)*240=60 16+60=76 (1/3)(4/4)-(1/4)(3/3)=(1/12) (1/3)*(1/12)=(1/36) (1/36)+(1/4)(9/9)=(10/36)=(5/18) a=76 b=(5/18)

OpenStudy (anonymous):

Thank you !